# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 3.2. Proof of Theorem 1
Finally, since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subsetneq\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ when $d\geq 2k-2$ (note that $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=2k-3<d$), the $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ is a non-trivial singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which means $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, by Lemma 10.
(iii) By assumption, $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-3,d\right\}=d$, that is to say, $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle={\mathbb{P}}^{d}$; hence the assertion follows.
(iv) For $(n,k)=(2,3)$, smoothness of all points in $\sigma_{3}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{2}(v_{d}({\mathbb{P}}^{2}))$ for $d\geq 4$ was already proved in [16, theorem 2.14]. This is included for completeness of the statement. ∎
@@
(iv) is the exception to the trichotomy in Theorem 1. Under the condition $(k,d,m,n)=(3,4,1,2)$ of (iv), the arithmetic deduced from the containment assumption (18) of moving tangents in the proof does not make any contradiction. The situation is similar in the other exceptional case to the trichotomy, $(k,d,m,n)=(4,3,2,3)$ (Part (iv) of Theorem 2).
@@ 4. Proof of main results
In this section, we prove Theorem 2 and Theorem 3. We will discuss the non-singularity result first and results for the singular loci next.
@@ 4.1. Generic smoothness
Assume $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={mk+k-1}$, and $k\leq\tbinom{m+d-1}{m}$. Fix $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ to be an $m$-plane, and take $a\in\sigma_{k}(v_{d}(L))$ to be a general point. Then, in the incidence $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$ with the first projection $p:I\rightarrow{\mathbb{P}}^{N}$ as (3) in §2, we have the following inclusion scheme-theoretically,
@@
Consider any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in p^{-1}(a)\subset I$. Let $I_{(m)}\subset I\subset{\mathbb{P}}^{N}\times(L)^{k}$ be the another incidence as in Lemma 9. Since $a\in\sigma_{k}(v_{d}(L))$ is general, it follows $a\notin\sigma_{k-1}(v_{d}(L))$ and $a\notin p(I_{(m)}\setminus I_{(m)}^{0})$ by Remark 7 (b). Since $\dim\sigma_{k}(v_{d}(L))={mk+k-1}$, the secant fiber of $I_{(m)}\rightarrow{\mathbb{P}}^{N}$ at $a$ also consists of finite points. So, by Lemma 9, we have $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I^{0}$. From Lemma 10, it is also true that $a\notin\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
We may write $a=\sum_{i=1}^{k}c_{i}x_{i}$ for some $c_{i}\in{\mathbb{C}}$ as regarding $a$ and $x_{i}=v_{d}(x_{i}^{\prime})$ as vectors in affine space ${\mathbb{C}}^{N+1}$. As in Remark 8 (a), we set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$. Then $a=\sum_{i=1}^{k}c_{i}y_{i}$, where it holds that $c_{i}\neq 0$ and $y_{i}^{\prime}\neq 0$ for all $1\leq i\leq k$. We may assume $x_{i}^{\prime}\in\left\{t_{0}\neq 0\right\}$ for all $i$.
@@
Since $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by $\mathsf{mono}[t,u]_{\leq d}$, and $a\in\langle\,v_{d}(L)\,\rangle$, it holds that
$0=\sum_{i=1}^{k}c_{i}\cdot\{(u_{1}\cdot\mathsf{mono}[t]_{\leq d-1})(x_{i}^{\prime})\}=\sum_{i=1}^{k}c_{i}\cdot x_{i,m+1}^{\prime}\cdot\{\mathsf{mono}[t]_{\leq d-1}(y_{i}^{\prime})\}~{},$
@@
where $\{Mono(pt)\}$ means the vector obtained by evaluating monomials in $Mono$ at the value of $pt$.
Applying Remark 7 (b) to $\sigma_{k}(v_{d}(L))$, we have
$\dim(\langle\,v_{d-1}(y_{1}^{\prime}),\dots,v_{d-1}(y_{k}^{\prime})\,\rangle)=k-1.$
@@
It means that $c_{i}\cdot x_{i,m+1}^{\prime}=0$, thus $x_{i,m+1}^{\prime}=0$ for all $1\leq i\leq k$. Similarly, we can obtain $x_{i,j}^{\prime}=0$ for all $1\leq i\leq k$ and $j>m$, which give the linear defining equations for $(L)^{k}$ in $({\mathbb{P}}^{n})^{k}$. Hence, scheme-theoretically, it follows that $x_{1}^{\prime},\dots,x_{k}^{\prime}\in L$. ∎
@@
We recall some known results on $k$-the secant variety and its incidence in terms of $k$-fold symmetric product of ${\mathbb{P}}^{n}$.
@@
    It is known that $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ is non-singular at $(x_{1}^{\prime},x_{2}^{\prime},\dots,x_{k}^{\prime})$ if $x_{i}^{\prime}\neq x_{j}^{\prime}$ whenever $i\neq j$. Thus, the subset of all distinct $k$-points of ${\mathbb{P}}^{n}$ is a smooth open subscheme of $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ (see e.g. [6, lemma 7.1.4]). Then we also consider the incidence variety in this setting as
    $\widetilde{I}\subset{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$
    , where $\widetilde{I}$ corresponds to $I$ in (3) under the natural map ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ and $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\subset{\mathbb{P}}^{N}$ is the first projection.
@@
    Assume $k(n+1)<\binom{n+d}{n}$. Then, from [9, theorem 1.1] we know that the projection $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is birational except $(k,d,n)=(9,6,2),(8,4,3),(9,3,5)$, because it is a dominant and generically injective morphism.
@@
For an $m$-plane $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$, we take the $m$-subsecant variety $Z=\sigma_{k}(v_{d}(L))$ of $Y=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. By assumptions of Theorems 2 and 3, we know that $\dim Y={nk+k-1}$ and $\dim Z={mk+k-1}\leq\dim\langle Z\rangle=\tbinom{m+d}{m}-1$ (i.e. $Z$ is non-defective) by Alexander-Hirschowitz theorem and $Z$ is also identifiable at its general point by [9, theorem 1.1]. Further, $(k,d,m)\neq(9,6,2),(8,4,3),(9,3,5)$ and note that $k\leq\tbinom{m+d-1}{m}$ for each $k$ of this range.
Let $a\in Z$ be a general point and consider $\widetilde{p}:\widetilde{I}\rightarrow Y$. By Lemma 25 and Remark 26 (b), we may assume that the scheme-theoretic fiber $\widetilde{p}^{-1}(a)$ is one point $\mathbf{x}=(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$ in $\widetilde{I}\cap({\mathbb{P}}^{N}\times\operatorname{Sym}^{k}(L))$ and $\mathbf{x}$ is a non-singular point in $\widetilde{I}$, because $\mathbf{x}$ is contained in a smooth Zariski open subset of $\widetilde{I}$ (i.e. a projective bundle over a smooth open base, see Remarks 7 (a) and 26 (a)).
Now, we restrict the projective morphism $\widetilde{p}:\widetilde{I}\rightarrow Y$ onto a non-empty affine open neighborhood $a\in W=\operatorname{Spec}A\subset Y$ and another open subset $\mathbf{x}\in U=\operatorname{Spec}B\subset\widetilde{I}$ and denote the injective ring homomorphism corresponding to $\widetilde{p}|_{U}:U\to W$ by $f:A\hookrightarrow B$ (let also $m_{a}$ (resp. $m_{\mathbf{x}}$) be the maximal ideal of $a$ in $A$ (resp. of $\mathbf{x}$ in $B$)). Note that we may take $U$ so that $\widetilde{p}|_{U}:U\to W$ is a finite morphism (cf. [18, chap. II, ex.3.22 (d)]).
Since $A/m_{a}\simeq B/m_{\mathbf{x}}\simeq{\mathbb{C}}$ and $\tilde{p}^{-1}(a)\simeq\operatorname{Spec}(B\otimes_{A}A/m_{a})\simeq\operatorname{Spec}(B/f(m_{a})B)$ is isomorphic to $\operatorname{Spec}{A/m_{a}}$, we have $f(m_{a})=m_{\mathbf{x}}$ in $B$ and $m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$ in the local ring $B_{m_{a}}$. We also know $A_{m_{a}}+m_{a}B_{m_{a}}/m_{a}B_{m_{a}}\simeq A_{m_{a}}/m_{a}A_{m_{a}}\simeq B_{m_{a}}/m_{a}B_{m_{a}}$, thus as $A_{m_{a}}$-module $A_{m_{a}}=B_{m_{a}}$ holds by Nakayama lemma. Note that $A_{m_{a}}=B_{m_{a}}$ is also true as a local ring. Finally, since we have $m_{a}A_{m_{a}}=m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$, we obtain the desired local ring isomorphism
$A_{m_{a}}=B_{m_{a}}\simeq(B_{m_{a}})_{m_{a}B_{m_{a}}}=(B_{m_{a}})_{m_{\mathbf{x}}B_{m_{a}}}\simeq B_{m_{\mathbf{x}}},$

### v3

@@ 3.2. Proof of Theorem 1
Finally, since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subsetneq\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ when $d\geq 2k-2$ (note that $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=2k-3<d$), the $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ is a non-trivial singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which means that $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, by Lemma 10.
(iii) By assumption, $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-3,d\right\}=d$, that is to say,
$\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle={\mathbb{P}}^{d};$
hence the assertion follows.
(iv) For $(n,k)=(2,3)$, smoothness of all points in $\sigma_{3}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{2}(v_{d}({\mathbb{P}}^{2}))$ for $d\geq 4$ was already proved in [Han18, Theorem 2.14]. This is included for completeness to the statement. ∎
@@
(iv) is the exception to the trichotomy in Theorem 1. Under the condition $(k,d,m,n)=(3,4,1,2)$ of (iv), the arithmetic deduced from the inclusion assumption (25) of moving tangents in the proof does not make any contradiction. The situation is similar in the other exceptional case to the trichotomy, $(k,d,m,n)=(4,3,2,3)$ (Part (iv) of Theorem 2).
@@ 4. Proof of main results
In this section, we prove Theorem 2 and Theorem 3. We will first discuss the non-singularity result and then the results for the singular loci.
@@ 4.1. Generic smoothness
Assume $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$ and $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={mk+k-1}$ (i.e., $v_{d}({\mathbb{P}}^{n})$ and $v_{d}({\mathbb{P}}^{m})$ are non-defective). Let $k\leq\tbinom{m+d-1}{m}$. Fix $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ to be an $m$-plane, and take $a\in\sigma_{k}(v_{d}(L))$ to be a general point. Then, in the incidence $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$ with the first projection $p:I\rightarrow{\mathbb{P}}^{N}$ as (4) in §2, we have the following inclusion scheme-theoretically,
@@
(i) Consider any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in p^{-1}(a)\subset I$. Let $I_{(m)}\subset{\mathbb{P}}^{N}\times(L)^{k}$ be the another incidence as in Lemma 9. Since $a\in\sigma_{k}(v_{d}(L))$ is general, it follows $a\notin\sigma_{k-1}(v_{d}(L))$ and $a\notin p(I_{(m)}\setminus I_{(m)}^{0})$ by Remark 7 (b). Since $\dim\sigma_{k}(v_{d}(L))={mk+k-1}$, the secant fiber of $I_{(m)}\rightarrow{\mathbb{P}}^{N}$ at $a$ also consists of finite points. So, by Lemma 9, we have $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I^{0}$. From Lemma 10, it is also true that $a\notin\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$. Thus, we may write $a=\sum_{i=1}^{k}c_{i}x_{i}$ for some $c_{i}\in{\mathbb{C}}$ as regarding $a$ and $x_{i}=v_{d}(x_{i}^{\prime})$ as vectors in the affine space ${\mathbb{C}}^{N+1}$, where $c_{i}\neq 0$ for all $1\leq i\leq k$.
@@
As in Remark 8, let us set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$. For $y_{i}=v_{d}(y_{i}^{\prime})$, the diagram (5) implies $a=\sum_{i=1}^{k}c_{i}y_{i}$, where $y_{i}^{\prime}\neq 0$ for $1\leq i\leq k$. For the affine open subset $V_{0}=\left\{t_{0}\neq 0\right\}\subset{\mathbb{P}}^{n}$, we may assume $x_{i}^{\prime}\in V_{0}$ for all $i$. Since $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized on $V_{0}$ by $\mathsf{mono}[t,u]_{\leq d}$, and $a\in\langle\,v_{d}(L)\,\rangle$, it holds that
$0=\sum_{i=1}^{k}c_{i}\cdot\{(u_{1}\cdot\mathsf{mono}[t]_{\leq d-1})(x_{i}^{\prime})\}=\sum_{i=1}^{k}c_{i}\cdot x_{i,m+1}^{\prime}\cdot\{\mathsf{mono}[t]_{\leq d-1}(y_{i}^{\prime})\},$
@@
where for $Mono=u_{1}\cdot\mathsf{mono}[t]_{\leq d-1},\mathsf{mono}[t]_{\leq d-1}$ and $pt=x_{i}^{\prime},y_{i}^{\prime}$, the symbol $\{Mono(pt)\}$ means the vector obtained by evaluating monomials in $Mono$ at the value of $pt$.
Since $k\leq\tbinom{m+d-1}{m}$, applying Remark 7 (b) to $\sigma_{k}(v_{d}(L))$, we may assume
(26) $\dim\langle\,v_{d-1}(y_{1}^{\prime}),\dots,v_{d-1}(y_{k}^{\prime})\,\rangle=k-1,$
@@
which gives $c_{i}\cdot x_{i,m+1}^{\prime}=0$, thus $x_{i,m+1}^{\prime}=0$ for all $1\leq i\leq k$ (more precisely, the linear independence of (26) means a $k$-minor of the corresponding matrix is non-zero, and $c_{i}\cdot x_{i,m+1}^{\prime}=0$ is obtained by multiplying the inverse of the $k\times k$ submatrix). Similarly, we can obtain $x_{i,j}^{\prime}=0$ for each $j>m$ and for all $1\leq i\leq k$, which gives the linear defining equations for $(L)^{k}$ in $({\mathbb{P}}^{n})^{k}$. Hence, $x_{1}^{\prime},\dots,x_{k}^{\prime}\in L$.
(ii) Let $U\subset({\mathbb{P}}^{n})^{k}$ be the open subset used in Remark 7, where $I^{0}$ is the ${\mathbb{P}}^{k-1}$-bundle over $U$. We define a morphism $\Phi:{\mathbb{P}}^{k-1}\times U\rightarrow{\mathbb{P}}^{N}\times U$ by
$\Phi((c_{1}:\dots:c_{k}),(x_{1}^{\prime},\dots,x_{k}^{\prime}))=\Bigl{(}\,\sum_{i=1}^{k}c_{i}v_{d}(x_{i}^{\prime}),\,(x_{1}^{\prime},\dots,x_{k}^{\prime})\Bigr{)}.$
Note that, by the linear independence of $v_{d}(x_{1}^{\prime}),\dots,v_{d}(x_{k}^{\prime})$ for $(x_{1}^{\prime},\dots,x_{k}^{\prime})\in U$, $\sum_{i=1}^{k}c_{i}v_{d}(x_{i}^{\prime})=\sum_{i=1}^{k}\tilde{c}_{i}v_{d}(x_{i}^{\prime})\in{\mathbb{P}}^{N}$ if and only if $(c_{1}:\dots:c_{k})=(\tilde{c}_{1}:\dots:\tilde{c}_{k})\in{\mathbb{P}}^{k-1}$. Then $\Phi({\mathbb{P}}^{k-1}\times U)=I^{0}$, and moreover, we have the isomorphism ${\mathbb{P}}^{k-1}\times U\simeq I^{0}$ under $\Phi$.
Let $U_{0}=U\cap(V_{0})^{k}\subset({\mathbb{P}}^{n})^{k}$, where $(V_{0})^{k}$ is an affine variety and its affine coordinates ring is $A={\mathbb{C}}[\{x_{i,j}^{\prime}\}]$. In addition, for each $k$-minor $\xi$ of the matrix whose $i$-th column consists of monomials of $m$ variables $x_{i,1}^{\prime},\dots,x_{i,m}^{\prime}$ of degrees $\leq d-1$, we set $(V_{0})^{k}_{\xi}=\left\{\xi\neq 0\right\}$, an open subset of $(V_{0})^{k}$ whose coordinates ring is $A_{\xi}$. Let $W\subset\left\{c_{1}\neq 0\right\}\subset{\mathbb{P}}^{k-1}$ be the affine open subset such that all the coordinates $c_{1},\dots,c_{k}$ are nonzero, where the coordinates ring of $W$ is ${\mathbb{C}}[c_{2},\dots,c_{k}]_{c_{2}\dotsm c_{k}}$ by regarding $c_{1}=1$.
We may assume $p^{-1}(a)\subset I^{0}\cap(\left\{a\right\}\times(V_{0})^{k})$. To consider the scheme-theoretic structure of $p^{-1}(a)$, for the composite morphism $\Phi_{1}=p\circ\Phi:{\mathbb{P}}^{k-1}\times U_{0}\rightarrow{\mathbb{P}}^{N}$, we take the fiber
$\Phi_{1}^{-1}(a)\subset{\mathbb{P}}^{k-1}\times U_{0}\subset{\mathbb{P}}^{k-1}\times(V_{0})^{k}.$
Since $a\notin\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, $\Phi_{1}^{-1}(a)\subset W\times(V_{0})^{k}$. Since $a$ is general in $\sigma_{k}(v_{d}(L))$, and by (26), $\Phi_{1}^{-1}(a)$ is contained in the union of affine open subsets $W\times(V_{0})^{k}_{\xi}$ with all $k$-minors $\xi$.
We take $F_{\xi}=\Phi_{1}^{-1}(a)\cap(W\times(V_{0})^{k}_{\xi})$ for each $\xi$, and consider the ideal $I(F_{\xi})$ in $A_{\xi}[c_{2},\dots,c_{k}]_{c_{2}\dotsm c_{k}}$, the affine coordinates ring of $W\times(V_{0})^{k}_{\xi}$. For $\beta=\binom{m+d}{m}-1$, we may write
$a=(1:a^{(1)}:\dots:a^{(\beta)}:0:\dots:0)\in\langle\,v_{d}(L)\,\rangle\subset{\mathbb{P}}^{N}$
with $a^{(1)},\dots,a^{(\beta)}\in{\mathbb{C}}$ and $a^{(\ell)}=0$ if $\ell>\beta$. Then the expression $a=\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})$ means that $a^{(\ell)}\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})^{(0)}-\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})^{(\ell)}\in{I(F_{\xi})}$ for $1\leq\ell\leq N$, where $v_{d}(x_{i}^{\prime})^{(\ell)}$ is the $\ell$-th coordinate of $v_{d}(x_{i}^{\prime})\in{\mathbb{P}}^{N}$. In particular, $\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})^{(\ell)}\in{I(F_{\xi})}$ for $\ell>\beta$. Using the discussion of (i), we have $x_{i,j}^{\prime}\in I(F_{\xi})$ for all $1\leq i\leq k$ and $j>m$, which means that $I(F_{\xi})$ contains the defining ideal of $({\mathbb{P}}^{k-1}\times(L)^{k})\cap(W\times(V_{0})^{k}_{\xi})$. Thus, scheme-theoretically, it follows $F_{\xi}\subset{\mathbb{P}}^{k-1}\times(U_{0}\cap(L)^{k})$ for any $\xi$, and hence $\Phi_{1}^{-1}(a)\subset{\mathbb{P}}^{k-1}\times(U_{0}\cap(L)^{k})$. Therefore, $p^{-1}(a)\subset\left\{a\right\}\times(L)^{k}$. ∎
@@
We recall some known results on the $k$-the secant variety and its incidence in terms of $k$-fold symmetric product of ${\mathbb{P}}^{n}$.
@@
    It is known that $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ is non-singular at $(x_{1}^{\prime},\dots,x_{k}^{\prime})$ if $x_{i}^{\prime}\neq x_{j}^{\prime}$ whenever $i\neq j$. Thus, the subset of all distinct $k$-points of ${\mathbb{P}}^{n}$ is a smooth open subscheme of $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ (see e.g. [BK05, Lemma 7.1.4]). Then we also consider the incidence variety in this setting as
    $\widetilde{I}\subset{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n}),$
    where $\widetilde{I}$ corresponds to $I$ in (4) under the natural map ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ and $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\subset{\mathbb{P}}^{N}$ is the first projection.
@@
    Assume $k(n+1)<\binom{n+d}{n}$. Then, we know from [COV17, Theorem 1.1] that the projection $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is birational except $(k,d,n)=(9,6,2),(8,4,3),(9,3,5)$, because it is a dominant and generically injective morphism.
@@
For an $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ with $m\geq 2$, we take the $m$-subsecant variety $Z=\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ of $Y=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. From [AH95], for $d\geq 3$, $Z$ does not fill $\langle\,Z\,\rangle$ and is secant defective if and only if $(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4)$. Thus, by the assumptions of Theorems 2 and 3, we know that
$\dim Y={nk+k-1},\quad\dim Z={mk+k-1}\leq\dim\langle\,Z\,\rangle=\tbinom{m+d}{m}-1,$
that is, $Y,Z$ are non-defective. In this case, $Z=\langle\,Z\,\rangle$ if and only if $k=\frac{\binom{m+d}{m}}{m+1}\in{\mathbb{N}}$. In particular, under the assumption $k<\mu$ of Theorem 2 (i), we have $Z\subsetneq\langle\,Z\,\rangle$.
If $Z\subsetneq\langle\,Z\,\rangle$, then since $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$ are excluded from Theorem 2 and (i) of Table 1, it follows from [COV17, Theorem 1.1] that $Z$ is generically identifiable. If $Z=\langle\,Z\,\rangle$, then $(k,d,m)=(5,3,3),(7,5,2)$ of (i) of Table 1 only occur, and in these cases, it follows from [GM17, Theorem 1] that $Z$ is generically identifiable.
Let $a\in Z$ be a general point and consider $\widetilde{p}:\widetilde{I}\rightarrow Y$. Note that $k\leq\tbinom{m+d-1}{m}$ for each $(k,d,m)$ of our range, which is an assumption of Lemma 25. Applying Lemma 25, Remark 26 (b), and the generic identifiability of $Z$, we may assume that the scheme-theoretic fiber $\widetilde{p}^{-1}(a)$ is one point $\mathbf{x}=(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in\widetilde{I}\cap({\mathbb{P}}^{N}\times\operatorname{Sym}^{k}(L))$ and $\mathbf{x}$ is a non-singular point in $\widetilde{I}$, because $\mathbf{x}$ is contained in a smooth Zariski open subset of $\widetilde{I}$ (i.e., a projective bundle over a smooth open base, see Remarks 7 (a) and 26 (a)).
Now, we restrict the projective morphism $\widetilde{p}:\widetilde{I}\rightarrow Y$ onto a non-empty affine open neighborhood $a\in W=\operatorname{Spec}A\subset Y$ and another open subset $\mathbf{x}\in U=\operatorname{Spec}B\subset\widetilde{I}$, and take the injective ring homomorphism $A\hookrightarrow B$ corresponding to $\widetilde{p}|_{U}:U\to W$. Also, let $m_{a}$ (resp. $m_{\mathbf{x}}$) be the maximal ideal of $a$ in $A$ (resp. of $\mathbf{x}$ in $B$). Note that we may take $U$ so that $\widetilde{p}|_{U}:U\to W$ is a finite morphism (cf. [H77, Chapter II, Ex.3.22 (d)]).
Since $A/m_{a}=B/m_{\mathbf{x}}\simeq{\mathbb{C}}$ and
$\widetilde{p}^{-1}(a)\simeq\operatorname{Spec}(B\otimes_{A}A/m_{a})\simeq\operatorname{Spec}(B/m_{a}B)$
is isomorphic to one simple point $\operatorname{Spec}{B/m_{\mathbf{x}}}$, we have $m_{a}B=m_{\mathbf{x}}$ in $B$. Let $B_{m_{a}}=B\otimes_{A}A_{m_{a}}$, whose member can be expressed as $b/s$ with $b\in B$ and $s\in A\setminus m_{a}$. We have $m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$ in $B_{m_{a}}$, and then,
$(A_{m_{a}}+m_{a}B_{m_{a}})/m_{a}B_{m_{a}}\simeq A_{m_{a}}/m_{a}A_{m_{a}}\simeq B_{m_{a}}/m_{a}B_{m_{a}},$
which implies $A_{m_{a}}+m_{a}B_{m_{a}}=B_{m_{a}}+m_{a}B_{m_{a}}=B_{m_{a}}$ as $A_{m_{a}}$-module. By Nakayama lemma (see e.g. [Matsu, Corollary of Theorem 2.2]), it follows $A_{m_{a}}=B_{m_{a}}$. In particular, $B_{m_{a}}$ is a local ring, whose maximal ideal is $m_{\mathbf{x}}B_{m_{a}}$. Thus, we have
$A_{m_{a}}=B_{m_{a}}=(B_{m_{a}})_{m_{\mathbf{x}}B_{m_{a}}}=B_{m_{\mathbf{x}}},$